OpenStudy (anonymous):

Identify the local maximum value for the function f (x) = x^6 - 6x^4 + 5, using the first derivative test.

OpenStudy (amistre64):

what do you get for the derivative?

OpenStudy (anonymous):

6x^5-24x^3

OpenStudy (anonymous):

similar to last one yes? take the derivative. factor. find the zeros. check the sign

OpenStudy (anonymous):

is that right?

OpenStudy (amistre64):

thats good

OpenStudy (anonymous):

it is cooked up to factor easily. and so easy to find the zeros yes it is right. factor now

OpenStudy (amistre64):

now to zero it out factor out what it has in common

OpenStudy (anonymous):

Ummm.. Im not sure how to do it but ill give it a shot.. :/ 6(x^5-4x^3). Is that right?

OpenStudy (anonymous):

Is this correct or am I wrong?

OpenStudy (anonymous):

can factor out a lot more than that. as in previous example

OpenStudy (anonymous):

i'll give you a hint. it is \[6x^3(x^2-4)\]

OpenStudy (anonymous):

okay. I got this far with your help.. I have no clue as to how to go to the next step

OpenStudy (anonymous):

set the derivative = 0 and solve

OpenStudy (anonymous):

you will get 3 zeros. name them

OpenStudy (anonymous):

\[6x^{3}\ +(x^{2}-4) = 0\]

OpenStudy (anonymous):

hello? pleae dont make me sad :(

OpenStudy (amistre64):

6x^5 -24x^3 ; factor out the 6 and the x^3 6x^3 (x^2 - 4) = 0 ; we can factor it again or just realize the +- 2 give the same results so when x = -2, 0, 2 we have critical points

OpenStudy (amistre64):

6x^3 (x-2) (x+2) when we plug in a -3 we get; -.-.- = - when we plug in a -1 we get; -.-.+ = + when we plug in a 1 we get; +.-.+ = - when we plug in a 3 we get; +.+.+ = + <.........-2..........0...........2.........> - + - + ride this like a roller coaster now, from left to right at -2 we go down then up, thats a minimum at 0 we go up then down; that a maximum at 2 we go down then up; that a minimum. http://www.wolframalpha.com/input/?i=x^6+-+6x^4+%2B+5